# MathMist: A Parallel Multilingual Benchmark Dataset for Mathematical Problem Solving and Reasoning

### v1

@@ 4.3 Prompt Techniques
We utilized established methods such as zero-shot prompting (Kuo and Chen, 2023) and Chain-of-Thought (CoT) (Wei et al., 2022) prompting. Beyond these standard techniques, we introduce a novel experimental setting termed Code-Switched CoT Prompting (CS-CoT). This approach adapts the sociolinguistic concept of code-switching (Aguilar et al., 2020; Hamed et al., 2025; Yan et al., 2025), the alternation between dissimilar languages in conversation. Specifically, this setting prompts the model to understand a mathematical question posed in one language while being instructed to generate the solution in another. To further probe the model’s analytical understanding, we present perturbed reasoning, an approach that evaluates the model’s proficiency to identify flawed steps within a given solution process. For details on the prompt templates used in this study, refer to Appendix A.1.
@@ 5.1 Cross-Lingual Performance in Zero-Shot and Chain-of-Thought Settings
The results in Table 3 provide a comprehensive view of how LLMs engage with human-annotated math word problems in both high- and low-resource languages under zero-shot and CoT prompting (see Appendix A.1). As shown in Table 6, Qwen, DeepSeek, and Mathstral demonstrate a substantial strength in proof-style problems but underperform in numerical and symbolic types under zero-shot conditions for both Bangla and English. In contrast, GPT-OSS-20B and Gemini-2.5-Flash-Lite reveal the contrasting tendency, achieving comparatively better results on numerical and symbolic problems. Notably, Gemini-2.5-Flash-Lite achieves the highest accuracy in the Bangla zero-shot setting but suffers a 15.77% drop in accuracy on the English zero-shot benchmark. Crucially, this deviation in behavior encourages a pointed and practical question: “Do Gemini models truly generalize mathematical problem solving beyond language?”. Subsequently, across all scenarios, GPT-OSS-20B attains the highest overall performance. CoT prompting significantly increases overall accuracy for most models, especially in English experiments. GPT-OSS-20B is an exception in that it demonstrates nearly consistent performance between zero-shot and CoT settings in both Bangla and English, while Gemini-2.5-Flash-Lite unexpectedly loses accuracy in Bangla when using CoT relative to Bangla zero-shot. Mathstral-7B shows poor performance, scoring below 30% in zero-shot settings and under 50% with CoT reasoning in both languages, equivalent to near-random baselines. These patterns reveal not just inter-model performance disparities but also a hidden sensitivity to problem type and language specificity.
@@
Figure 2: Mathematical reasoning performance across seven typologically diverse languages. Models with broad multilingual pretraining achieve higher accuracy, while specialized models show greater cross-lingual variability.
@@ 5.2 Cross-lingual Performance Patterns on MCQs: Why Deterministic Answer Spaces Amplify Model Accuracy?
We stratified performance by language typology and resource level, utilizing an arbitrary sample per language (50 numerical, 50 symbolic, 50 proof problems); all problems were identical across languages. The scale and diversity of multilingual pretraining greatly impacted accuracy and consistency, as illustrated in Figure 2. GPT-OSS-20B, which shows the highest accuracy ($\mu\approx 74.2\%$, $\sigma\approx 2.9\%$), while Mathstral-7B wilted ($\mu\approx 23.0\%$, $\sigma\approx 5.7\%$). Mid-size models like Qwen-8B ($\mu\approx 46.2\%$) and DeepSeek R1-7B ($\mu\approx 40.9\%$) performed moderately. Gemini-2.5-Flash-Lite matched mid-range accuracy ($\mu\approx 47.8\%$) but with higher variance ($\sigma\approx 9.9\%$). The biases associated with each language clearly exhibit the consequences of both pretraining and fine-tuning. GPT-OSS-20B model evidently favors high-resource languages like English and French, while the Gemini performs adequately well in Bangla. Qwen shows good performance in both French and Turkish. In contrast, DeepSeek R1 exhibits noteworthy weaknesses with Kazakh. Mathstral performs poorly in English and French, but toils even more with typologically distant languages such as Kazakh and Turkish. Overall, these patterns show that broad multilingual pretraining is more useful than limited specialization for cross-linguistic mathematical reasoning. Inconsistent representation and limited fine-tuning can create fragile, language-biased results, highlighting the need for multilingual-aware fine-tuning, targeted data augmentation, and lightweight adapters for underrepresented language families.
@@
A breakdown by linguistic code-switching, specifically the mismatch between the question and the CoT language, shows a clear performance penalty across all evaluated LLMs, revealing that linguistic alignment is critical for mathematical reasoning generalization as depicted in Figure 3. GPT-OSS-20B demonstrates the highest resilience, with a smaller decline in accuracy (from 77.58% to 73.29%), showcasing its strong generalization capability. Conversely, the smaller models, DeepSeek R1 7B and Mathstral 7B, showed considerably larger declines in performance, indicating that they have constrained generalization capabilities and depend laboriously on training language and cross-lingual consistency.
@@
This indicates that while larger parameters provide a performance boost, effective cross-lingual reasoning and linguistic grounding are paramount to minimize accuracy drops from linguistic transitions in complex tasks. The accuracy breakdown by category is shown in Table 7. Table 8 to 15 illustrates the language proportions in the chain-of-thought solutions when instructed to use Bangla (see Appendix A.1).
@@
We evaluated 150 sampled problems with an equal distribution of Numerical, Symbolic, and Proof types. The experiment utilized Strategy 3 (discussed in 3.3) to test the models. Subsequently, Qwen-8B excels in binary classification tasks but struggles significantly with complex error identification, implying that high-level classification isn’t a reliable measure of analytical depth. For instance, models such as Mathstral 7B and DeepSeek R1 7B demonstrate a drastic divergence, indicating their reasoning pathways are insufficient for deep error analysis. This distinction indicates a failure to generalize from simple detection to complex reasoning. Cross-lingual understanding plays a pivotal role, as demonstrated by GPT-OSS 20B outperforming others in Bangla error identification, suggesting that targeted instruction tuning for step-by-step logical analysis can be more useful than sheer parameter scale. We evaluated LLM outputs against perturbation-generated ground truth, automatically judged by an LLM and manually validated across all 300 problems. The consistent performance decline in Bangla versus English reveals significant language effects, indicating limited generalization in low-resource languages—a crucial barrier to cross-lingual mathematical reasoning.
@@ 6 Discussion
Our experiments revealed that in LLM mathematical problem-solving, zero-shot performance favored symbolic and proofs, while CoT prompting performed better in numerical problems. A significant failure mode was code-switching reasoning, specifically with Bangla CoT, which repeatedly led to linguistically broken mathematical outputs containing mixed tokens. Furthermore, we observed several concerning behaviors, including models producing over 4,600 tokens without a solution before concluding with "Hmm, I’m stuck", self-imposed limits ("I think I’ve reached the limit of my understanding"), and deterministic backtracking that induced hallucinations and infinite recursive loops (see Appendix A.2). Finally, Gemini underperformed in the English zero-shot setting compared to its Bangla performance and against open-source models, suggesting it may rely on pattern-matching rather than deep mathematical reasoning.
Models #Param Bangla Error Detection Error Identification Accuracy (TP rate) F1 Score GPT-OSS 20B 78.00% 58.67% $\downarrow$ 73.95% Qwen 3 8B 83.33% 56.00% $\downarrow$ 71.83% DeepSeek R1 7B 66.67% 24.00% $\downarrow$ 38.72% Mathstral 7B 78.00% 20.00% $\downarrow$ 33.34% Models #Param English Error Detection Error Identification Accuracy (TP rate) F1 Score GPT-OSS 20B 79.33% 61.33% $\downarrow$ 76.03% Qwen 3 8B 88.00% 68.67% $\downarrow$ 81.42% DeepSeek R1 7B 72.00% 25.33% $\downarrow$ 40.43% Mathstral 7B 81.33% 24.00% $\downarrow$ 38.72%
@@ 7 Conclusion
While MathMist offers a robust pipeline for cross-lingual math benchmarking, it has several limitations that suggest future research directions. First, our exclusive use of zero-shot prompting techniques, Chain-of-Thought (CoT), Code-Switching Reasoning, and Perturbed Reasoning with frozen models could be improved by supervised fine-tuning. Secondly, incorporating a one-shot, few-shot example or using advanced strategies such as Atom of Thoughts (AoT) (Teng et al., 2025) or Program of Thoughts (PoT) (Chen et al., 2022) can enhance reasoning traces. Thirdly, the dataset’s quality and coverage could be remarkably improved by replacing synthetic translations with human validation and expanding the curriculum from secondary to higher secondary school mathematics. Finally, our analysis could be deepened by benchmarking against a broader range of state-of-the-art proprietary models, such as GPT-5, Gemini-2.5 Pro, or Claude-4, to provide a more comprehensive evaluation of cross-lingual math-solving capabilities.
@@
All data used in MathMist were sourced from publicly available secondary school mathematics textbooks authorized by the Government of Bangladesh, ensuring that no personally identifiable or sensitive information is included. The dataset focuses solely on academic content and adheres to fair use and research ethics guidelines. The resource is intended strictly for educational and research purposes to advance equitable, transparent, and linguistically inclusive AI development in mathematical reasoning.
@@ A.2 Qualitative Examples
A.3 Breakdown of Quantitative Analysis
The tables detail an evaluation of LLM mathematical reasoning across English and Bangla, focusing on the efficacy of Chain-of-Thought (CoT) prompting and the problem of linguistic interference. The data reveals that CoT generally boosts performance (Table 6), especially for smaller models, but high-performing models show minimal or negative gains in Bangla, indicating linguistic brittleness. This pervasive linguistic inconsistency (Table 7) identifies a major challenge to cross-lingual robustness and instruction fidelity. Moreover, the language composition tables (8) expose widespread code-switching reasoning during Bangla CoT: smaller models (DeepSeek R1, Mathstral) exhibit "Mixed" language outputs, injecting English, Cyrillic, etc. (Table 10).
@@ A.3 Breakdown of Quantitative Analysis
Models #Param Bangla Zero-Shot CoT Numerical Symbolic Proof Overall Numerical Symbolic Proof Overall Qwen 8B 54.92 47.24 64.10 55.36 73.67 ($\uparrow$18.75) 68.28 ($\uparrow$21.04) 71.47 ($\uparrow$7.37) 72.11 ($\uparrow$16.75) GPT-OSS 20B 79.00 73.45 62.50 74.33 78.29 ($\downarrow$0.71) 71.72 ($\downarrow$1.73) 66.03 ($\uparrow$3.53) 75.22 ($\uparrow$0.89) DeepSeek R1 7B 32.38 35.17 67.31 40.48 61.09 ($\uparrow$28.71) 54.83 ($\uparrow$19.66) 66.03 ($\downarrow$1.28) 60.90 ($\uparrow$20.42) Mathstral 7B 24.08 25.52 56.73 31.42 35.47 ($\uparrow$11.39) 37.59 ($\uparrow$12.07) 59.94 ($\uparrow$3.21) 41.18 ($\uparrow$9.76) Gemini-2.5-Flash-Lite $\approx$2B 79.95 71.72 65.06 75.09 73.07 ($\downarrow$6.88) 70.69 ($\downarrow$1.03) 65.38 ($\downarrow$0.68) 71.38 ($\downarrow$3.71) Models #Param English Zero-Shot CoT Numerical Symbolic Proof Overall Numerical Symbolic Proof Overall Qwen 8B 37.49 42.76 61.22 43.67 78.05 ($\uparrow$40.56) 72.76 ($\uparrow$30.00) 72.44 ($\uparrow$11.22) 75.78 ($\uparrow$32.11) GPT-OSS 20B 79.12 75.52 68.59 76.12 80.90 ($\uparrow$1.78) 75.52 ($\uparrow$0.00) 70.51 ($\uparrow$1.92) 77.58 ($\uparrow$1.46) DeepSeek R1 7B 41.28 40.34 65.06 46.23 66.79 ($\uparrow$25.51) 60.34 ($\uparrow$19.99) 63.78 ($\downarrow$1.28) 64.84 ($\uparrow$18.61) Mathstral 7B 21.23 27.93 53.85 29.62 40.33 ($\uparrow$19.10) 38.97 ($\uparrow$11.04) 64.42 ($\uparrow$10.57) 45.26 ($\uparrow$15.64) Gemini-2.5-Flash-Lite $\approx$2B 64.18 59.31 59.29 62.15 74.14 ($\uparrow$9.96) 71.38 ($\uparrow$12.07) 65.06 ($\uparrow$5.77) 71.63 ($\uparrow$9.48)

### v2

@@ 4.3 Prompt Techniques
We utilized established methods such as zero-shot prompting (Kuo and Chen, 2023) and Chain-of-Thought (CoT) (Wei et al., 2022) prompting. Beyond these standard techniques, we introduce a novel experimental setting termed Code-Switched CoT Prompting (CS-CoT).
This approach adapts the sociolinguistic concept of code-switching (Aguilar et al., 2020; Hamed et al., 2025; Yan et al., 2025), which refers to the alternation between dissimilar languages in conversation. While recent studies have begun to explore cross-lingual reasoning, they often treat code-switching as an input-side feature (Chai et al., 2025) or rely on the model’s English-centric capabilities to bridge the reasoning gap. For instance, (Son et al., 2025b) conducts reasoning primarily in English while maintaining only the final mathematical semantics in the target language.
Figure 2: Mathematical reasoning performance across 13 typologically diverse languages. Models with broad multilingual pretraining achieve higher accuracy, while specialized models show greater cross-lingual variability.
In contrast, our setting prompts the model to understand a mathematical question posed in one language while being instructed to generate the solution in another. To further probe the model’s analytical understanding, we present perturbed reasoning, an approach that evaluates the model’s proficiency to identify flawed steps within a given solution process. For details on the prompt templates used in this study, refer to Appendix A.1.
@@ 5.1 Cross-Lingual Performance in Zero-Shot and Chain-of-Thought Settings
The results in Table 4 provide a comprehensive view of how LLMs engage with human-annotated math word problems in both high- and low-resource languages under zero-shot and CoT prompting (see Appendix A.1). As shown in Table 9, Qwen, DeepSeek, and Mathstral demonstrate a substantial strength in proof-style problems but underperform in numerical and symbolic types under zero-shot conditions for both Bangla and English (see Appendix A.6). In contrast, GPT-OSS-20B and Gemini-2.5-Flash-Lite reveal the contrasting tendency, achieving comparatively better results on numerical and symbolic problems. Notably, Gemini-2.5-Flash-Lite achieves the highest accuracy in the Bangla zero-shot setting but suffers a 15.77% drop in accuracy on the English zero-shot benchmark. Crucially, this deviation in behavior encourages a pointed and practical question: “Do Gemini models truly generalize mathematical problem solving beyond language?”. Subsequently, across all scenarios, GPT-OSS-20B attains the highest overall performance. CoT prompting significantly increases overall accuracy for most models, especially in English experiments. GPT-OSS-20B is an exception in that it demonstrates nearly consistent performance between zero-shot and CoT settings in both Bangla and English, while Gemini-2.5-Flash-Lite unexpectedly loses accuracy in Bangla when using CoT relative to Bangla zero-shot. Mathstral-7B shows poor performance, scoring below 30% in zero-shot settings and under 50% with CoT reasoning in both languages, equivalent to near-random baselines. These patterns reveal not just inter-model performance disparities but also a hidden sensitivity to problem type and language specificity. Additionally, we conducted an error bar analysis of GPT-OSS 20B across five runs on the Bangla dataset, reporting a mean accuracy of 75.96% (standard deviation of 1.16; 95% confidence interval [73.69, 78.23]). Category-level variances were as follows: Numerical (77.64 $\pm$ 1.03), Symbolic (74.54 $\pm$ 1.79), and Proof (72.86 $\pm$ 5.07). These findings indicate that MathMist effectively captures variations in reasoning performance across various mathematical task types.
@@ 5.2 Cross-lingual Performance Patterns on MCQs: Why Deterministic Answer Spaces Amplify Model Accuracy?
The scale and diversity of multilingual pretraining greatly impacted accuracy and consistency, as illustrated in Figure 2. GPT-OSS-20B, which shows the highest accuracy ($\mu\approx 72.11\%$, $\sigma\approx 3.83\%$), while Mathstral-7B wilted ($\mu\approx 21.58\%$, $\sigma\approx 6.83\%$). Mid-size models like Qwen-8B ($\mu\approx 43.81\%$) and DeepSeek R1-7B ($\mu\approx 38.10\%$) performed moderately, but Qwen-8B shows the highest variance ($\sigma\approx 13.49\%$). Gemini-2.5-Flash-Lite matched mid-range accuracy ($\mu\approx 53.66\%$). The biases associated with each language clearly exhibit the consequences of both pretraining and fine-tuning. GPT-OSS-20B model evidently favors high-resource languages like English and French, while the Gemini performs adequately well in Bangla and French. Qwen shows good performance in both Persian and Bangla. In contrast, DeepSeek R1 exhibits noteworthy weaknesses with Hausa. Mathstral performs poorly in Hausa, but toils even more with typologically distant languages such as Amharic. Overall, these patterns show that broad multilingual pretraining is more useful than limited specialization for cross-linguistic mathematical reasoning. Inconsistent representation and limited fine-tuning can create fragile, language-biased results, highlighting the need for multilingual-aware fine-tuning, targeted data augmentation, and lightweight adapters for underrepresented language families. Overall category-level results for the full dataset are shown in Table 11.
@@
A breakdown by linguistic code-switching, specifically the mismatch between the question and the CoT language, shows a clear performance penalty across all evaluated LLMs, revealing that linguistic alignment is critical for mathematical reasoning generalization as depicted in Figure 3. GPT-OSS-20B demonstrates the highest resilience, with a smaller decline in accuracy (from 77.58% to 73.29%), showcasing its strong generalization capability.
@@
Conversely, the smaller models, DeepSeek R1 7B and Mathstral 7B, showed considerably larger declines in performance, indicating that they have constrained generalization capabilities and depend laboriously on training language and cross-lingual consistency. This indicates that while larger parameters provide a performance boost, effective cross-lingual reasoning and linguistic grounding are paramount to minimize accuracy drops from linguistic transitions in complex tasks. The accuracy breakdown by category is shown in Table 10. Table 12 to 19 illustrates the language proportions in the chain-of-thought solutions when instructed to use Bangla (see Appendix A.1).
@@
Strategy Model English Bangla Error Detection Error Identification Error Detection Error Identification $\sigma_{1}$ Qwen-8B 79.10% 55.85% $\downarrow$ 79.79% 51.90% $\downarrow$ GPT-OSS-20B 67.82% 51.90% $\downarrow$ 67.20% 54.39% $\downarrow$ DeepSeek R1-7B 54.33% 24.78% $\downarrow$ 51.14% 20.21% $\downarrow$ Mathstral-7B 76.12% 19.65% $\downarrow$ 71.70% 13.08% $\downarrow$ $\sigma_{2}$ Qwen-8B 71.07% 56.12% $\downarrow$ 68.44% 49.07% $\downarrow$ GPT-OSS-20B 47.54% 39.86% $\downarrow$ 44.64% 36.40% $\downarrow$ DeepSeek R1-7B 39.72% 15.16% $\downarrow$ 40.76% 15.22% $\downarrow$ Mathstral-7B 70.17% 18.69% $\downarrow$ 70.17% 13.84% $\downarrow$ $\sigma_{3}$ Qwen-8B 86.99% 62.98% $\downarrow$ 83.88% 56.96% $\downarrow$ GPT-OSS-20B 78.06% 62.63% $\downarrow$ 74.46% 58.06% $\downarrow$ DeepSeek R1-7B 57.16% 24.08% $\downarrow$ 52.87% 18.34% $\downarrow$ Mathstral-7B 77.72% 20.90% $\downarrow$ 75.02% 14.88% $\downarrow$
Strategy $\sigma_{3}$ outperforms others because it combines multiple fallacies, leading to cascading inconsistencies that create stronger error signals. Unlike $\sigma_{1}$ and $\sigma_{2}$, where errors are limited to one or two steps, $\sigma_{3}$ introduces multiple interacting mistakes that create obvious contradictions throughout the solution. These widespread inconsistencies are easier for language models to detect across different languages. Subsequently, Qwen-8B excels in binary classification tasks but struggles significantly with complex error identification, implying that high-level classification isn’t a reliable measure of analytical depth. For instance, models such as Mathstral 7B and DeepSeek R1 7B demonstrate a drastic divergence, indicating their reasoning pathways are insufficient for deep error analysis. This distinction indicates a failure to generalize from simple detection to complex reasoning. Cross-lingual understanding plays a pivotal role, as demonstrated by GPT-OSS 20B outperforming others in Bangla error identification, suggesting that targeted instruction tuning for step-by-step logical analysis can be more useful than sheer parameter scale. We evaluated LLM outputs against perturbation-generated ground truth, with assessments conducted using an LLM-subject matter expert (SME) pipeline 4.2. The consistent performance decline in Bangla versus English reveals significant language effects, indicating limited generalization in low-resource languages—a crucial barrier to cross-lingual mathematical reasoning.
@@ 6 Discussion
Our experiments revealed that in LLM mathematical problem-solving, zero-shot performance favored symbolic and proofs, while CoT prompting performed better in numerical problems. Moreover, scaling Qwen-3 (0.6B to 14B) improves English reasoning accuracy from 26.85% to 45.95% and narrows the English-Bangla performance gap from 7% to 1% (see Appendix A.5). This scale-driven alignment almost doubles the accuracy rate of Bangla logical proofs to 70%, demonstrating that larger models effectively reduce language-specific discrepancies in complex reasoning. A significant failure mode was code-switching reasoning, specifically with Bangla CoT, which repeatedly led to linguistically broken mathematical outputs containing mixed tokens. Furthermore, we observed several concerning behaviors, including models producing over 4,600 tokens without a solution before concluding with "Hmm, I’m stuck", self-imposed limits ("I think I’ve reached the limit of my understanding"), and deterministic backtracking that induced hallucinations and infinite recursive loops (see Appendix A.2). Finally, Gemini underperformed in the English zero-shot setting compared to its Bangla performance and against open-source models, suggesting it may rely on pattern-matching rather than deep mathematical reasoning.
@@ 7 Conclusion
While MathMist offers a robust pipeline for cross-lingual math benchmarking, it has several limitations that suggest future research directions. First, our exclusive use of zero-shot prompting techniques, Chain-of-Thought (CoT), Code-Switching Reasoning, and Perturbed Reasoning with frozen models could be improved by supervised fine-tuning. Secondly, incorporating a one-shot, few-shot example or using advanced strategies such as Atom of Thoughts (AoT) (Teng et al., 2025) or Program of Thoughts (PoT) (Chen et al., 2022) can enhance reasoning traces. Thirdly, the dataset’s coverage could be remarkably improved by expanding the curriculum from secondary to higher secondary school mathematics. Finally, our analysis could be deepened by benchmarking against a broader range of state-of-the-art proprietary models, such as GPT-5, Gemini-2.5 Pro, or Claude-4, to provide a more comprehensive evaluation of cross-lingual math-solving capabilities.
@@
All data used in MathMist were sourced from publicly available secondary school mathematics textbooks authorized by the People’s Republic of Bangladesh, ensuring that no personally identifiable or sensitive information is included. The dataset focuses solely on academic content and adheres to fair use and research ethics guidelines. The resource is intended strictly for educational and research purposes to advance equitable, transparent, and linguistically inclusive AI development in mathematical reasoning.
@@ A.2 Qualitative Examples
A.3 Language Coverage in the MathMist Benchmark
This appendix provides detailed statistics of the languages included in the MathMist benchmark.
[…]
Table 7: Detailed Statistics of the Languages in MathMist Benchmark
Table 7 summarizes the linguistic diversity of the dataset, reporting each language’s ISO code, language family, writing script, and an approximate estimate of the number of native and second-language speakers worldwide. The selected languages encompass various language families and writing systems, including Indo-European, Afro-Asiatic, Niger-Congo, Turkic, and Uralic. They represent both left-to-right and right-to-left scripts, as well as Semitic and Bantu morphologies. Together, these languages represent an estimated total speaker reach of approximately 3.15 billion, highlighting the broad global coverage of MathMist. This diversity enables a systematic evaluation of multilingual mathematical reasoning across typologically and script-wise heterogeneous languages.
A.4 LLM-as-a-Judge: Evaluation Protocol and Scoring Rubric
We conducted a human review with an SME (subject-matter expert), which is standard for mathematical evaluation requiring domain precision. Table 3 shows a 0 to 100 human scoring rubric that covers correctness, alignment of reasoning, and clarity. The SME evaluated and ensured uniformity without inter-annotator variability. Additionally, the SME performed back-translation checks using Google Translate and DeepL to verify semantic fidelity across languages. The LLM-as-a-Judge achieved a Human Consistency Score ranging from 91.39/100 to 94.8/100 across non-MCQ experiments. This approach aligns closely with expert reasoning. Minor deductions were mostly due to explanations that were correct but slightly under-specified. This analysis reinforces the reliability of the judge’s decisions and their alignment with expert human understanding.
A.5 Analysis of Scaling Effects on the Qwen Model Family
We conducted an ablation study with the Qwen 3 model family, ranging from 0.6B to 14B parameters, to analyze scaling effects on mathematical reasoning in English and Bangla. Our results confirm a strong positive correlation between parameter size and performance, particularly in English, where accuracy improved from 26.85% at 0.6B to 45.95% at 14B. Larger models significantly reduce the performance gap between English and Bangla. The smallest model (0.6 billion parameters) showed a 7% advantage for English over Bangla, while the largest model (14 billion parameters) resulted in nearly equal performance, with only a 1% difference. Furthermore, the ability to solve logical proofs improved with model size. In Bangla, success rates nearly doubled from 38% with the smaller model to 70% with the larger model. We have included a detailed breakdown of these results in the table 8 below.
@@ A.5 Analysis of Scaling Effects on the Qwen Model Family
Model Bangla Accuracy (%) English Accuracy (%) Numerical Symbolic Proof Overall Numerical Symbolic Proof Overall Qwen 0.6B 12.93 18.97 38.78 19.72 16.61 21.72 59.29 26.85 Qwen 1.7B 12.34 ($\downarrow$0.59) 23.10 ($\uparrow$4.13) 52.88 ($\uparrow$14.10) 23.25 ($\uparrow$3.53) 26.69 ($\uparrow$10.08) 21.03 ($\downarrow$0.69) 54.17 ($\downarrow$5.12) 31.49 ($\uparrow$4.64) Qwen 4B 64.53 ($\uparrow$51.60) 61.72 ($\uparrow$42.75) 56.41 ($\uparrow$17.63) 62.21 ($\uparrow$42.49) 32.62 ($\uparrow$16.01) 26.90 ($\uparrow$5.18) 60.26 ($\uparrow$0.97) 37.44 ($\uparrow$10.59) Qwen 8B 54.92 ($\uparrow$41.99) 47.24 ($\uparrow$28.27) 64.10 ($\uparrow$25.32) 55.36 ($\uparrow$35.64) 37.49 ($\uparrow$20.88) 42.76 ($\uparrow$21.04) 61.22 ($\uparrow$1.93) 43.67 ($\uparrow$16.82) Qwen 14B 36.30 ($\uparrow$23.37) 42.76 ($\uparrow$23.79) 70.19 ($\uparrow$31.41) 44.91 ($\uparrow$25.19) 34.40 ($\uparrow$17.79) 47.59 ($\uparrow$25.87) 75.64 ($\uparrow$16.35) 45.95 ($\uparrow$19.10)
Table 8: Scaling effects of the Qwen-3 model family (0.6B–14B) on mathematical reasoning in Bangla and English. Relative changes ($\uparrow$ / $\downarrow$) indicate performance differences with respect to the smallest model (Qwen-3:0.6B) within the same language and category. Green highlights the highest accuracy achieved in each column. Results demonstrate a strong positive correlation between model size and reasoning performance, particularly for proof-based tasks and overall accuracy.
We marked unexpectedly high numerical reasoning spikes in the Qwen-3:4B Bangla checkpoint, presumably due to specific data mixture variance in that release. But the general trend confirms that larger models offer stable cross-lingual alignment.
A.6 Breakdown of Quantitative Analysis
The tables detail an evaluation of LLM mathematical reasoning across English and Bangla, focusing on the efficacy of Chain-of-Thought (CoT) prompting and the problem of linguistic interference. The data reveals that CoT generally boosts performance (Table 9), especially for smaller models, but high-performing models show minimal or negative gains in Bangla, indicating linguistic brittleness.
Models #Param Bangla Zero-Shot CoT Numerical Symbolic Proof Overall Numerical Symbolic Proof Overall Qwen 8B 54.92 47.24 64.10 55.36 73.67 ($\uparrow$18.75) 68.28 ($\uparrow$21.04) 71.47 ($\uparrow$7.37) 72.11 ($\uparrow$16.75) GPT-OSS 20B 79.00 73.45 62.50 74.33 78.29 ($\downarrow$0.71) 71.72 ($\downarrow$1.73) 66.03 ($\uparrow$3.53) 75.22 ($\uparrow$0.89) DeepSeek R1 7B 32.38 35.17 67.31 40.48 61.09 ($\uparrow$28.71) 54.83 ($\uparrow$19.66) 66.03 ($\downarrow$1.28) 60.90 ($\uparrow$20.42) Mathstral 7B 24.08 25.52 56.73 31.42 35.47 ($\uparrow$11.39) 37.59 ($\uparrow$12.07) 59.94 ($\uparrow$3.21) 41.18 ($\uparrow$9.76) Gemini-2.5-Flash-Lite N/A 79.95 71.72 65.06 75.09 73.07 ($\downarrow$6.88) 70.69 ($\downarrow$1.03) 65.38 ($\downarrow$0.68) 71.38 ($\downarrow$3.71) Models #Param English Zero-Shot CoT Numerical Symbolic Proof Overall Numerical Symbolic Proof Overall Qwen 8B 37.49 42.76 61.22 43.67 78.05 ($\uparrow$40.56) 72.76 ($\uparrow$30.00) 72.44 ($\uparrow$11.22) 75.78 ($\uparrow$32.11) GPT-OSS 20B 79.12 75.52 68.59 76.12 80.90 ($\uparrow$1.78) 75.52 ($\uparrow$0.00) 70.51 ($\uparrow$1.92) 77.58 ($\uparrow$1.46) DeepSeek R1 7B 41.28 40.34 65.06 46.23 66.79 ($\uparrow$25.51) 60.34 ($\uparrow$19.99) 63.78 ($\downarrow$1.28) 64.84 ($\uparrow$18.61) Mathstral 7B 21.23 27.93 53.85 29.62 40.33 ($\uparrow$19.10) 38.97 ($\uparrow$11.04) 64.42 ($\uparrow$10.57) 45.26 ($\uparrow$15.64) Gemini-2.5-Flash-Lite N/A 64.18 59.31 59.29 62.15 74.14 ($\uparrow$9.96) 71.38 ($\uparrow$12.07) 65.06 ($\uparrow$5.77) 71.63 ($\uparrow$9.48)